A 400.0-N sign hangs from the end of a uniform strut. The strut is 4.0m long and weighs 600.0 N. The strut is supported by a hinge at the wall and by a cable whose other end is tied to the wall at a point 3.0m above the left end of the strut. Find the tension in supporting cable and the force of the hinge on the strut

1
Expert's answer
2021-12-09T09:15:16-0500

The tension in the supporting cable


"\sum\tau_B=0\to600\cdot2+400\cdot4-4T\cdot(3/5)=0\to T=1167\ (N)" . Answer


The force of the hinge on the strut.


"F_h=T\cdot(4/5)=1167\cdot(4/5)=933\ (N)"


"\sum\tau_A =0\to600\cdot2-F_v\cdot4=0\to F_v=300\ (N)"


"F=\sqrt{933^+300^2}=980\ (N)" . Answer